Consider the table.

Shirts Pants
Year Price Quantity Price Quantity
2013 $5.00 2000 $15.00 3000
2014 $8.00 1900 $20.00 2500

The table contains prices and output for a two good economy. Nominal and real GDP in 2013 are both $55,000. Use the information in the table to answer the questions. What is nominal GDP in 2014?

What is real GDP in 2014? Use 2013 as the base year

Real output from 2013 to 2014 has

increased.

stayed the same.

experienced inflation.

decreased.
Answer #1

Nominal gdp in 2014= 8*1900 + 20*2500= 65200

Real gdp in 2014= 5*1900 + 15*2500= 47000

Real gdp has DECREASED